Answer the following problems.

1. Find the values of t for which the area on the right tail of the t-distribution is 0.05 and the number of degrees of freedom is equal to:

a. 15 b. 28 c. 100


2. Find the 99th percentile of the t-distribution with 18 degrees of freedom.


3. Find the 90th percentile of the t-distribution if the sample size is 25.


1
Expert's answer
2021-06-27T17:33:33-0400

1.

a. "t=1.75305"


b. "t=1.701131"


c. "t=1.660234"


2.

99th percentile "(\alpha=0.01)"

"t=2.55238"


3.

90th percentile "(\alpha=0.10), df=n-1=25-1=24"

"t=1.317836"